# LMM-Assisted Breast Cancer Treatment Target Segmentation with Consistency Embedding

### v2

@@ Abstract
Recent advancements in Artificial Intelligence (AI) have profoundly influenced medical fields, by providing tools to reduce clinical workloads. However, most AI models are constrained to execute unimodal tasks, in stark contrast to the comprehensive approaches utilized by medical professionals. To address this, here we present RO-LMM, a multi-purpose large multimodal model (LMM) tailored for the field of radiation oncology. This model covers series of tasks within clinical workflow, adept at clinical report summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation. In particular, to perform consecutive clinical tasks, we further present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the capability of handling clean inputs, and transform this concept into LMM-driven segmentation framework as Consistency Embedding Segmentation (CESEG). Experimental results on multi-centre cohorts demonstrate our RO-LMM’s promising performance for multiple clinical tasks with generalization capabilities.
@@
Figure 1: RO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology. The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation. All the patient information and any potential personally identifiable information are de-identifiied, and the clinical data usage is approved by IRB.
**footnotetext: These authors contributed equally to this work\$$\dagger$\$$\dagger$\$footnotetext: Corresponding authors
1 Introduction
@@
Recently, the emergence of a new generation of AI models, known as foundation models, marks a significant departure from previous paradigms [34]. These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [3, 10, 12, 27, 31, 25].
@@
These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision-making. Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [21]. Specifically, in contrast to these unimodal AIs, generalist medical AIs can encompass a holistic understanding of clinical workflows, which can receive and reason a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [41, 46, 39, 50, 34].
Inspired by this paradigm shift in medical AIs, here we present RO-LMM, a prototype of a multi-purpose medical large multimodal model (LMM), tailored to support the clinical workflow in radiation oncology. RO-LMM is adept at performing a variety of clinical tasks: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes. Additionally, it is capable of (2) proposing appropriate treatment plans from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed treatment plans. This multifaceted functionality demonstrates a significant alignment with the expertise of clinical professionals. To the best of our knowledge, our prototype for assisting in the radiation oncology workflow is the first work. As for an initial trial, we opted for a phased approach, starting with high prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging, i.e., breast cancer.
@@
In the course of the aforementioned sequential tasks, the inevitable accumulation of errors in generations arises on both treatment plan suggestion and target volume segmentation task. To address this issue, we firstly explore Noisy Embedding Fine-Tuning (NEFTune) [19] which involves injecting uniform noise into embeddings during the training for our targeted tasks. To further enhance the model’s applicability, we introduce a novel consistency regularization method, bolstered by our pioneering Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs. Expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG). These advancements collectively contribute to a marked improvement in the model’s generalization capabilities in both internal and external validation. Our contributions can be summarized as:
@@
  • We propose a comprehensive framework, denoted as RO-LMM, wherein LMM assists the entire workflow of radiation oncology for breast cancer treatment.
  • We explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG.
@@
  • Through experiments on both internal and external datasets on real clinical breast cancer data, we demonstrated that RO-LMM outperforms baseline methods.
2 Related Works
2.1 Instruction Fine-tuning
@@
Instruction fine-tuning has been emerged as a pivotal technique for augmenting model responsiveness in the field of natural language processing (NLP). The simplicity and effectiveness of this approach have been introduced in numerous works, with a notable scale-up facilitated by advancements of large language models (LLMs) such like GPT-3 [2], ChatGPT [36], and GPT-4 [37]. Notably, a pioneering contribution, Self-Instruct [47], involves fine-tuning of foundation models using instruction-output pairs generated from InstructGPT [38]. This methodology, along with similar approaches, has led to diverse language model variants, including Alpaca [42], Vicuna [4], Dolly [8], and a series of LLaMA [43, 44], by exhibiting promising performance across a wide range of tasks. In the medical domain, Chat-Doctor [52], Med-Alpaca [14], PMC-LLaMA [49], and Asclepius [24] have been fine-tuned for clinical question-answering (QA) task. Although these models demonstrate robust performance for diverse QA benchmarks, there is a notable gap in their applicability for the practical clinical workflow in a specific domain, such as radiation oncology, which serves as our targeted focus.
2.2 Stabilizing LLM Fine-tuning with Noise
@@
Figure 2: Training schematics of the multi-purpose RO-LMM for three different tasks. (a) RO-LMM-S for clinical note summarization. (b) RO-LMM-P++ for radiation treatment plan suggestion. (c) RO-LMM-SEG++ for plan-guided target volume segmentation.
2.3 Language-driven Image Segmentation
@@
Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [26], open-vocabulary segmentation [11, 51, 29], referring segmentation [48], and reasoning segmentation [25]. Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable. For instance, LViT [28] and ConTEXTualNet [18] introduce text-driven chest X-ray radiography segmentation.
@@
Target volume segmentation in the field of radiation oncology is more challenging, due to its intrinsic need for considerations of the clinical aspects beyond the image, such as overall cancer stage, treatment aim, pathological findings, and so on [5, 6, 15]. Recent research on clinical context-aware breast cancer radiotherapy delineation has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce [35]. Unlike previous approaches, we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
3 Methods
@@
In this section, we provide a detailed methodology for our proposed approach, which is designed for sequential text generation tasks, including summarization and suggestions, as well as text-driven image segmentation, as illustrated in Figure 2. In Sec 3.1, we describe text-related fine-tuning approaches that yield report summarization and plan suggestion, respectively. In Sec 3.2, we introduce a novel segmentation approach that involves injecting noise and regularizing consistency to the text embedding for improving robustness of plan-guided 3D target volume segmentation.
3.1 Consistency Embedding Fine-tuning for Clinical LMM
@@
To realize the multi-purpose LMM with expertise in clinical report summarization and radiation treatment plan suggestion, we conduct instruction fine-tuning for LLaMA2 [44]. Considering the intended objective of each task, we adopt separate strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert). We train a summary expert using collected notes. During inference, the summary expert receives raw clinical reports, just like in the training scenario. However, for the plan expert, there is a discrepancy between the training and inference scenarios. In other words, we opt to use the training set made up of collected notes instead of generated notes, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure 1. In contrast to training, our model takes the generated notes as input for inference. To enhance robustness for the noisy input, Noisy Embedding Fine-Tuning (NEFTune) [19] which inject uniform noise into embedding is also an effective naive solution to handle noisy inputs in this task. However, a crucial consideration arises from the nature of the generated notes, which may lie closer to clean inputs (collected notes) or deviate towards noisy inputs. To address this, it is essential to train the model to handle both clean and noisy inputs. To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++†$\dagger$†$\dagger$†$\dagger$+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively. More details are as follows.
@@
Revisiting Noise Embedding Fine-Tuning. NEFTune has been recently introduced to enhance the performance of instruction fine-tuning. This approach involves injecting a random noise vector into embeddings during the training process as follows:
@@
${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}={\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\boldsymbol{\epsilon,\quad\epsilon}\sim\mathcal{U}(-1,1)$ (1)
@@
Consistency Embedding Fine-Tuning. To bolster the model’s capability in handling both noisy and clean inputs for treatment plan suggestion task, we incorporate a regularization loss to encourage consistency as follows:
@@
where $\theta^{-}$ represents the model with stopped gradient, and $d(\cdot,\cdot)$ is used to quantify the discrepancy between $f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ and $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$. The function $\mathcal{T}$ serves as either the identity function or performs detokenization to generate sentences. Our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs. The metric function $d$ can be any capable of measuring distance in a specific space. However, solely minimizing distance in the same embedding space might weaken the robustness effect and pose limitations on enforcing semantic similarity. In our approach, to capture textual similarity between sentences, we calculate the distance in the feature space of SentenceBERT [40]. i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [40], $\mathcal{T}$ is set to detokenization operator. We provide algorithmic descriptions of CEFT in Supplementary Material.
3.2 LMM-assisted 3D Target Volume Segmentation
For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LMM for context-aware 3D segmentation framework introduced in [35]. However, as the LMM-assisted 3D segmentation model gets the generated treatment plan as text conditions which are inevitably noisy from the previous summarization step, we introduce noise augmentation and consistency regularization modules for the target volume segmentation network, namely, Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG).
Noise Embedding Segmentation (NESEG). Firstly, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, and denote it as NESEG. The loss function for NESEG can be formulated as:
@@ 3.2 LMM-assisted 3D Target Volume Segmentation
where $\Theta$ = [$\psi,\phi$], ${\mathbf{x}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D CT scan, $B$ denotes batch size, $H$, $W$, and $S$ correspond to height, width, and slice of the CT scan, ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ is the perturbed text embedding, ${\mathbf{y}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D ground truth segmentation mask, $g_{\psi}(\cdot)$ represents the multimodal segmentation model parameterized by $\psi$, $f_{\theta^{*}}$ denotes a frozen LLM, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is the noise injected plan embeddings, and $\mathbf{z}_{\phi}$ is the learnable text prompt parametreized by $\phi$. We denote this model as RO-LMM-SEG+.
Consistency Embedding Segmentation (CESEG). We further adapt the consistency regularization module for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan, and denote it as CESEG. Apart from the original concept of CEFTune, we modified the consistency regularization module for the multimodal segmentation task, which combines CEFTune and the text prompt tuning. Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is regularized with that of clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. The loss function for CESEG is formulated as:
@@
where, $\phi^{-}$ represents the learnable text prompts with stopped gradient, $d(\cdot,\cdot)$ denotes the cosine similarity, and ${\mathbf{y}^{\text{emb}}_{\text{txt}}}=f_{\theta^{*}}({\mathbf{x}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi^{-}})\in\mathbb{R}^{B\times 1\times C}$ is the text embedding for the clean plan ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. The algorithm for CESEG is provided in Supplementary Material.
@@
The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 3. We employ the 3D Residual U-Net [7] as for the backbone structure within the image module, as shown in (a). In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input. In the multimodal alignment module, as illustrated as (c), we adopt both self-attention and cross-attention mechanisms in an interactive manner (image-to-text and text-to-image) for each $l$-th skip-connection layer, by following the concept of promptable segmentation borrowed from Segment Anything Model (SAM) [23]). For realizing multimodal alignment within the 3D segmentation framework, we adapt the idea of light-weight text prompt tuning to transfer the linguistic capability of LLM, by following the concepts introduced in [35, 22, 53]. Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module. Further details of each module are deferred to Supplementary Material.
4 Experiments
4.1 Dataset

### v3

# End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding

@@ Abstract
Recent advances in AI foundation models have significant potential for lightening the clinical workload by mimicking the comprehensive and multi-faceted approaches used by medical professionals. In the field of radiation oncology, the integration of multiple modalities holds great importance, so the opportunity of foundational model is abundant. Inspired by this, here we present RO-LMM, a multi-purpose, comprehensive large multimodal model (LMM) tailored for the field of radiation oncology. This model effectively manages a series of tasks within the clinical workflow, including clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation by leveraging the capabilities of LMM. In particular, to perform consecutive clinical tasks without error accumulation, we present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the consistency of handling clean inputs. We further extend this concept to LMM-driven segmentation framework, leading to a novel Consistency Embedding Segmentation (CESEG) techniques. Experimental results including multi-centre validation confirm that our RO-LMM with CEFTune and CESEG results in promising performance for multiple clinical tasks with generalization capabilities.
Index Terms:
Large multimodal model, Breast cancer, Clinical report, Radiotherapy target volume, Segmentation
I Introduction
[…]
@@ I Introduction
figureRO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology. The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and 3D target volume segmentation.
@@
Recently, the emergence of a new generation of AI models inspired by the large language models (LLMs), known as foundation models, marks a significant departure from previous paradigms [1]. These models are characterized by their massive scale and versatility, which stems from self-supervised training on a diverse array of data. These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [2, 3, 4, 5, 6, 7].
@@
These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision-making. Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [8]. Specifically, in contrast to these unimodal AIs, generalist medical AIs that synergistically combine foundation models can encompass a holistic understanding of clinical workflows, which can receive a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [9, 10, 11, 1, 12]. By understanding diverse data types and their interrelationships, multimodal AIs would facilitate more accurate diagnoses, personalized treatment development, and a reduction in medical errors by providing a comprehensive view of patient data.
In the field of radiation oncology, which the main focus of this article, the integration of multiple modalities holds great importance so that it is one of the most important clinical fields to evaluate the potential of foundation models. Accordingly, here we introduce RO-LMM, a prototype large multimodal model (LMM) specifically designed to support the clinical workflow in radiation oncology. In particular, this work significantly extends our related prior work called LLMSeg[13], which focused on multimodal segmentation. More specifically, RO-LMM enlarges the LLMSeg’s scope by tackling a broader range of clinical tasks in radiation oncology: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes. Additionally, it is capable of (2) proposing appropriate treatment plans from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed treatment plans. This multifaceted functionality of RO-LLM demonstrates a significant advances toward the expertise of clinical professionals.
@@
In training LLM to perform these sequential tasks from treatment plan suggestion to target volume segmentation, we found that there exist potentials for error accumulation from each tasks which can potential leads to significant end-to-end performance loss. Accordingly, another important contribution of this work is the employment and extension of a Noisy Embedding Fine-Tuning (NEFTune) [14] which involves injecting uniform noise into embeddings during the training for each targeted task. More specifically, to further enhance the model’s applicability, we develop a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs. Furthermore, by expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in novel Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG). These advancements prevent from error propagation from each subsequent task and collectively contribute to a marked improvement in the end-to-end model’s generalization capabilities in both internal and external validation. As for the proof of the concept study, our RO-LLM framework was adopted for breast cancer study, one of the high-prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging. Our contributions can be summarized as:
@@
  • We propose a comprehensive framework, dubbed as RO-LMM, wherein LMM assists the broad workflow of radiation oncology for breast cancer treatment. To the best of our knowledge, our prototype is the first to assist in the comprehensive workflow of radiation oncology.
  • To prevent from potential error accumulation throughout consecutive clinical tasks such as clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation, we explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG that significantly enhance the robustness of our method.
@@
  • Through experiments on various validation settings with real clinical data of breast cancer patients, we demonstrated that RO-LMM outperforms baseline methods.
II Related Works
II-A Instruction Fine-tuning in LLM
@@
In the field of LLM, the instruction fine-tuning has been emerged as a pivotal technique for augmenting model responsiveness. The simplicity and effectiveness of this approach have been introduced in numerous works, with a notable scale-up facilitated by advancements of large language models (LLMs) such like GPT-3 [15], ChatGPT [16], and GPT-4 [17]. Notably, a pioneering contribution, Self-Instruct [18], involves fine-tuning of foundation models using instruction-output pairs generated from InstructGPT [19]. This methodology, along with similar approaches, has led to diverse language model variants, including Alpaca [20], Vicuna [21], Dolly [22], and LLaMA2 [23], by exhibiting promising performance across a wide range of tasks. In the medical domain, Chat-Doctor [24], Med-Alpaca [25], PMC-LLaMA [26], and Asclepius [27] have been fine-tuned for clinical question-answering (QA) task. Although these models demonstrate robust performance for diverse QA benchmarks, there is a notable gap in their applicability for the practical clinical workflow in a specific domain, such as radiation oncology, which serves as our targeted focus.
II-B Stabilizing LLM Fine-tuning with Noise
@@
Figure 1: Schematics of RO-LMM training for three different tasks. (a) RO-LMM-S for clinical note summarization. (b) RO-LMM-P++ for radiation treatment plan suggestion. (c) RO-LMM-SEG++ for plan-guided target volume segmentation.
II-C Language-driven Image Segmentation
@@
Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [32], open-vocabulary segmentation [33, 34], referring segmentation [35], and reasoning segmentation [7]. Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable. For instance, LViT [36] and ConTEXTualNet [37] introduce text-driven chest X-ray radiography segmentation.
@@
Target volume segmentation in the field of radiation oncology is more challenging, due to its intrinsic need for considerations of the clinical aspects beyond the image, such as overall cancer stage, treatment aim, pathological findings, and so on [38, 39, 40]. Recent research on clinical context-aware breast cancer radiotherapy delineation [13] has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce. By extending [13], we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
III Methods
@@
In this section, we provide a detailed description of our proposed approach designed for sequential text generation tasks, including summarization and suggestions, as well as text-driven image segmentation, whose robustness is improved by consistency embedding finetuning. The overall framework is illustrated in Figure 1.
III-A Consistency Embedding Fine-tuning for Clinical LMM
To realize the multi-purpose LMM with expertise in clinical report summarization and radiation treatment plan suggestion, we conduct instruction fine-tuning for LLaMA2 [23]. Considering the nuanced differences in the intended objective of each task, we adopt separate training strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert).
@@ III-A Consistency Embedding Fine-tuning for Clinical LMM
Specifically, we train a summary expert using collected raw clinical report and summary notes. During inference, the summary expert receives raw clinical reports, just like in the training scenario. However, for the plan expert, there is a discrepancy between the training and inference scenarios. In other words, we use the training set made up of collected summary notes instead of generated notes from the summary expert, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure I. However, at the inference phase, our model takes the generated notes from the trained summary expert. To deal with the input domain differences for the training and inference time, Noisy Embedding Fine-Tuning (NEFTune) [14] which inject uniform noise into embedding could be an effective naive solution to handle noisy inputs in this task. However, a crucial consideration arises from the nature of the generated notes, since some of them may lie closer to clean inputs (collected notes) and the others deviate towards noisy inputs. To address this, it is essential to train the model to handle both clean and noisy inputs. To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++111‘+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively. More details are as follows.
@@
Consistency Embedding Fine-Tuning. To enhance the performance of instruction fine-tuning, NEFTune  [14] involves injecting a random noise vector into embeddings during the training process as follows:
@@
${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}={\mathbf{x}^{\text{emb}}_{\text{txt}}}+(\alpha/\sqrt{LC})\bm{\epsilon,\quad\epsilon}\sim\mathcal{U}(-1,1)$ (1)
@@
To bolster the model’s capability in handling both noisy and clean inputs for treatment plan suggestion task, we incorporate a regularization loss to encourage consistency as follows:
@@
where $\theta^{-}$ represents the model with stopped gradient, and $d(\cdot,\cdot)$ is used to quantify the discrepancy between $f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ and $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$. The metric function $d$ can be any capable of measuring distance in a specific space. The function $\mathcal{T}$ performs detokenization to generate sentences.
By introducing the consistency in the detokenized domain through $\mathcal{T}$, our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs. Since solely minimizing distance in the same embedding space may weaken the robustness and pose limitations on enforcing semantic similarity, to capture textual similarity between sentences, another contribution of our work is to calculate the distance in the feature space of SentenceBERT [41]. i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [41].
III-B LMM-assisted 3D Target Volume Segmentation
For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LLM for context-aware 3D segmentation introduced in LLMSeg[13]. LLMSeg demonstrated significant performance improvement in multimodal radiotherapy target segmentation compared to unimodal models. Additionally, it showed the ability to align target contours with provided clinical text information. However, in the extended clinical workflow proposed here, the LMM-assisted segmentation model receives noisy text descriptions of the generated treatment plan from the previous summarization step. To address this noise, we introduce two novel modules: Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG). These modules enhance the target volume segmentation network by handling noise and promoting consistency. The details are as follows.
Noise Embedding Segmentation (NESEG). Firstly, we introduce the noise augmentation technique, dubbed as NESEG, for the target volume segmentation task. We refer to the corresponding RO-LLM as RO-LMM-SEG+. Specifically, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, leading to the follow loss function:
@@ III-B LMM-assisted 3D Target Volume Segmentation
where $\Theta$ = [$\psi,\phi$], ${\mathbf{x}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D CT scan, $B$ denotes batch size, $H$, $W$, and $S$ correspond to height, width, and slice of the CT scan, ${\mathbf{\tilde{y}}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times 1\times C}$ is the perturbed text embedding, ${\mathbf{y}_{\text{img}}}\in\mathbb{R}^{B\times HWS}$ is the 3D ground truth segmentation mask, $g_{\psi}(\cdot)$ represents the multimodal segmentation model parameterized by $\psi$, $f_{\theta^{*}}$ denotes a frozen LLM, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is the noise injected plan embeddings, and $\mathbf{z}_{\phi}$ is the learnable text prompt parametreized by $\phi$.
Consistency Embedding Segmentation (CESEG). We further adapt the consistency regularization module, dubbed as CESEG, for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan. The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 2.
Apart from the original concept of CEFTune, CESEG modifies the consistency regularization module for the multimodal segmentation task, by combining CEFTune and the text prompt tuning. Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is also regularized with respect to the clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. Specifically, the loss function for CESEG is formulated as:
@@
where, $\phi^{-}$ represents the learnable text prompts with stopped gradient, $d(\cdot,\cdot)$ denotes the cosine similarity, and ${\mathbf{y}^{\text{emb}}_{\text{txt}}}=f_{\theta^{*}}({\mathbf{x}^{\text{emb}}_{\text{txt}}};\mathbf{z}_{\phi^{-}})\in\mathbb{R}^{B\times 1\times C}$ is the text embedding for the clean plan ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$.
@@
We employ the 3D Residual U-Net [42] as for the backbone structure within the image module, as shown in (a). In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input. In the multimodal alignment module, as illustrated as (c), we adopt both self-attention and cross-attention mechanisms in an interactive manner (image-to-text and text-to-image) for each $l$-th skip-connection layer, by following the concept of promptable segmentation borrowed from Segment Anything Model (SAM) [43]). For realizing multimodal alignment within the 3D segmentation framework, we adapt the idea of light-weight text prompt tuning to transfer the linguistic capability of LLM, by following the concepts introduced in [13, 44, 45]. Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module.
IV Experiments
IV-A Dataset